OpenStudy (kinggeorge):

[SOLVED] Prove that \[\sum_{j=0}^n (-1)^j \binom{n}{j} = 0\]For all \(n > 0\).

OpenStudy (kinggeorge):

I feel like I'm missing some basic trick.

OpenStudy (amistre64):

induction?

OpenStudy (kinggeorge):

Supposedly, but I'm not sure how to set it up properly.

OpenStudy (amistre64):

I dont see any values to avoid perse

OpenStudy (amistre64):

lets create a base case

OpenStudy (amistre64):

is it true for n=1?

OpenStudy (kinggeorge):

\[\binom{1}{0} - \binom{1}{1}=1-1=0\]Check.

OpenStudy (amistre64):

good, how about 2 and 3 for good measure?

OpenStudy (amistre64):

and isnt this just the pascal rows? with alternating signs ....

OpenStudy (kinggeorge):

\[\binom{2}{0} - \binom{2}{1} + \binom{2}{1}=1-2+1=0\]\[\binom{3}{0} - \binom{3}{1} + \binom{3}{2} - \binom{3}{3}=1-3+3-1=0\]Check.

OpenStudy (kinggeorge):

So we can assume it's true for some \(n=k\). My issue is then extending that to \(k+1\).

OpenStudy (amistre64):

1-4+6-4+1 = 8-8 = 0

OpenStudy (amistre64):

well, lets write up the k proposal wo we have something to wirk with

OpenStudy (kinggeorge):

I see a combinatorics argument we could make depending on the fact that n is odd or even.

OpenStudy (amistre64):

i think pascals is proved inductively too isnt it

OpenStudy (kinggeorge):

As for the \(k\) proposal...\[\sum_{j-0}^k (-1)^j \binom{k}{j} = 0\]Show that \[\sum_{j-0}^{k+1} (-1)^j \binom{k+1}{j} = 0\]

OpenStudy (amistre64):

yeah, that second one is familiar to the proof for the pascals ....

OpenStudy (amistre64):

id have to google it to iron out the details

OpenStudy (kinggeorge):

I see the correct relation to Pascal's rule now. Let's see where that takes us.

OpenStudy (kinggeorge):

\[\sum_{j-0}^{k+1} (-1)^j \binom{k+1}{j} = \sum_{j-0}^{k} (-1)^j \left[\binom{k}{j} + \binom{k}{j-1}\right]\]So we can distribute, and then by our hypothesis, we know that it's true since we're using \(k\) instead of \(k+1\).

OpenStudy (anonymous):

How about Expanding \( (1-x)^n\) and then substituting \(x=1\)?

OpenStudy (kinggeorge):

I suppose that might just be a better/easier way to do it.

OpenStudy (anonymous):

You suppose? It is actually :)

OpenStudy (kinggeorge):

Maybe I just like the more complicated ways better. :)

OpenStudy (anonymous):

Good luck :)